OpenStudy (anonymous):

Solve. 4(a + 2) = 8 + 4a all real numbers no solution –2 –1

OpenStudy (catlover5925):

ok if you distribute 4 you get --> 4a + 8 = 8 + 4a then subtract 4a from both sides and you end up with 8=8...so I think that is all real #'s

OpenStudy (dmndlife24):

After distributing...you'll find that both sides equal each other, 4a+2=4a+2

OpenStudy (dmndlife24):

Sorry 4a + 8 = 4a + 8

OpenStudy (dmndlife24):

This is a true statement which implies all real numbers

OpenStudy (anonymous):

thank you both :)

OpenStudy (dmndlife24):

Your welcome :D

OpenStudy (catlover5925):

awww I didn't get medaled :(

OpenStudy (dmndlife24):

Here u go...

OpenStudy (dmndlife24):

You actually answered first anyway

OpenStudy (catlover5925):

aww thanks!!